The thermal decomposition of dimethyl ether (CH3)2O(g) -> CH4(g) + H2(g) + CO(g) Is to be carried out in an isothermal 2.0 L lab reactor at 600 oC. The reactor is charged with pure dimethyl ether at a pressure of 349 torr. After about two hours, the reactor pressure is 840 torr. Calculate the percentage of the dimethyl ether has decomposed.
Answer #1

CH3)2O(g) -> CH4(g) + H2(g) + CO(g)
349 0 0 0 (initial)
349-x x x x (at later timr)

total pressure at later time = 349-x+x+x+x = 349 + 2x
840 = 349 + 2x
x = 245.5 torr

percent decomposition = x*100/initial p
= 245.5 * 100 / (349)
= 70.3 %
Answer: 70.3 %